Adam pushes a 47 kg boat across a frozen lake. The coefficient of kinetic friction between the boat and the ice is 0.13 and the boat acceleration at 0.6 m/s^2. Calculate the magnitude of Adams force in the boat.
1
Expert's answer
2019-10-14T09:15:38-0400
"ma=F-F_{friction}=F-\mu mg"

The magnitude of Adams force on the boat:


"F=ma+\mu mg=m(a+\mu g)"

"F=47(0.6+(0.13)(9.8))=88\ N"